Problem 9-25 Direct Materials and Direct Labor Variances; Computations from Incomplete Data [LO9-4, LO9-5]

Sharp Company manufactures a product for which the following standards have been set:

Standard Quantity
or Hours
Standard Price
or Rate
Standard
Cost
Direct materials 3 feet $ 5 per foot $ 15
Direct labor ? hours ? per hour ?

During March, the company purchased direct materials at a cost of $59,040, all of which were used in the production of 3,000 units of product. In addition, 4,600 direct labor-hours were worked on the product during the month. The cost of this labor time was $34,500. The following variances have been computed for the month:

Materials quantity variance $ 4,200 U
Labor spending variance $ 3,900

U

Labor efficiency variance $ 680

U

Required:

1. For direct materials:

a. Compute the actual cost per foot of materials for March.

b. Compute the price variance and the spending variance.

2. For direct labor:

a. Compute the standard direct labor rate per hour.

b. Compute the standard hours allowed for the month’s production.

c. Compute the standard hours allowed per unit of product.
Answer #1

Solution 1a (Sharp Company):
Material quantity variance = $4,200 U
(SQ - AQ)*SP = -$4,200
(3000*3 - AQ) * $5 = - $4,200
Actual quantity = 9840 Foot
Actual cost of purchase = $59,040
actual cost per foot = $59040 / 9840 = $6.00 per foot

Solution 1b:
Material price variance = (SP - AP) * AQ = ($5 - $6) * 9840 = $9,840 U
Material spending variance = MPV + MUV = $9840 U + $4200 U = $14,040 U

Solution 2a:
Actual rate of labor per hour = $34,500 / 4600 = $7.50 per hour
Labor rate variance = Labor spending variance - Labor efficiency variance = $3,900 U - $680 U = $3,220 U
(SR - AR) * AH = - $3220
(SR - $7.50) * 4600 = - $3220
SR = $6.80 per hour


Solution 2b:
Labor efficiency variance = (SH - AH) * SR

(SH - 4600) * $6.80 = - $680
Standard hours = 4500 hours


solution 2c:
Standard hours allowed per unit = 4500 / 3000 = 1.50 hour per unit.